OpenStudy (anonymous):

given a parallelogram ABCD (diagonals from A to C and D to B. where the diagonals cross is E.), find the following: AE= 3x+2y EC= 25 AB= 2x+9y Cd= 5x+y-5. how do i find x and y

OpenStudy (ranga):



OpenStudy (ranga):

In a parallelogram: the opposite sides are of equal length. Therefore, 2x+9y = 5x+y-5 the diagonals bisect each other. Therefore, AE = EC or 3x+2y = 25. You have two equations and two unknowns. So you can solve for x and y.

OpenStudy (anonymous):

every time i work out that problem it never equals out. help.

OpenStudy (ranga):

2x+9y = 5x+y-5 5x + y - 5 - 2x - 9y = 0 3x - 8y = 5 ----- (1) 3x + 2y = 25 ------ (2) subtract (1) from (2) 10y = 20 y = 2. put y = 2 in (2) 3x + (2)(2) = 25 3x + 4 = 25 3x = 21 x = 7 x = 7; y = 2.

OpenStudy (anonymous):

also for that same picture but for these DF=2x=y CF=15 AB=3x+2y DA=5x-3y+3 (f is where the diagonals cross.)

OpenStudy (ranga):

Are you sure the vertices have the letters A,B,C,D counter-clockwise? If not, you can click on the Draw button below the reply box and draw and label the diagram.

OpenStudy (anonymous):

no it is ABCD starting at the left top corner for this one.

OpenStudy (ranga):

You are given two adjacent sides of the parallelogram as (3x+2y) and (5x+3y+3)? And one half of each diagonal as (2x+y) and 15?

OpenStudy (ranga):

It would make the question easier to understand if you'd draw it, label the vertices and write down the length of each segment near it.

OpenStudy (ranga):

You can also take a pic of the problem and upload it. If it online, you can take a screen shot with "print screen" on your keyboard in Windows and upload the screenshot.